OpenStudy (anonymous):

Find the value of x in the logarithm log2 x = -3. A. 1/3 B. 3/2 C. 1/8 D. 8

OpenStudy (anonymous):

1/3

OpenStudy (harsimran_hs4):

log2 x = -3 x = 2^(-3) x = 1/8